Let f : A → B be a function.


1. Show that for the identity function iA on A we have f ◦ iA = f.


2. Show that for the identity function iB on B we have iB ◦ f = f.





1
Expert's answer
2021-12-07T11:42:04-0500

"f:\ A\rightarrow B"

1. "\forall a \in A:\quad (f\circ I_A)(a)=f\big(I_A(a)\big)"

Since "a\in A" , it follows that "I_A(a)=a" and "(f\circ I_A)(a)=f\big(I_A(a)\big)=f(a)"

So, "f\circ I_A=f" .


2. "\forall a \in A:\quad (I_B\circ f)(a)=I_B\big(f(a)\big)"

Since "f(a)\in B" , it follows that "I_B\big(f(a)\big)=f(a)" and "(I_B\circ f)(a)=f(a)"

So, "I_B\circ f=f" .